# Evaluation Agent: Efficient and Promptable Evaluation Framework for Visual Generative Models

### v1

@@ Abstract
Fan Zhang1∗, Shulin Tian2∗, Ziqi Huang2∗, Yu Qiao1🖂, Ziwei Liu2🖂 1Shanghai Artificial Intelligence Laboratory, 2S-Lab, Nanyang Technological University zhangfan2@pjlab.org.cn, {shulin002, ziqi002}@ntu.edu.sg https://vchitect.github.io/Evaluation-Agent-project/
00footnotetext: equal contributions. 🖂corresponding authors.   Code is available
@@ 2.1 Visual Generation and Evaluation
Visual generative models have gained significant attention in recent years. However, unlike perception tasks, which have clear evaluation metrics such as accuracy, evaluating visual generative tasks is more challenging due to the absence of a definitive “ground truth” or single correct answer. Metrics such as FID Heusel et al. (2017) and FVD Unterthiner et al. (2018) are commonly used to measure the distance between generated samples and reference datasets. Recent benchmarks, including T2I-CompBench Huang et al. (2023), VBench Huang et al. (2024a), and EvalCrafter Liu et al. (2023), provide multi-dimensional evaluations tailored to specific model capabilities. However, whether relying on metrics like FID Heusel et al. (2017) and FVD Unterthiner et al. (2018) or benchmarks like VBench Huang et al. (2024a), these evaluations are often time-intensive and limited in scope.
@@ 3.1 Preliminaries: Evaluation of Visual Generative Models
[…]
@@ A.2 Execuation Stage
Appendix B Experiment Implementation Details
For T2V tasks, we validate our evaluation approach by comparing the consistency of our Evaluation Agent with VBench’s original evaluation scheme across multiple dimensions, with respect to evaluation time, sample count, and final assessment results on four open-source T2V models. For VBench, we selected 15 evaluation dimensions, including Subject Consistency, Background Consistency, and Motion Smoothness, among others. Details could be referred to in Table 3.1. We used VBench’s original scheme to sample and evaluate four models on these dimensions, recording generation time, evaluation time, and results. We categorized $5$ tiers based on the sample-level results from VBench: Very High, High, Moderate, Low, and Very Low, based on the scores density distributions. For dimensions scored by proportions, like Dynamic Degree, we used scores from $42$ models on VBench’s leaderboard for tiering. To ensure fairness, we restricted the evaluation tools and prompts to those in VBench’s existing metrics. Our Evaluation Agent assessed each dimension by answering targeted questions, invoking relevant tools, and selecting suitable prompts, guided by detailed definitions and score-tiering information to interpret results and provide an overall performance assessment for each model.
Appendix C Open-Ended User Query Dataset
C.1 Building an Open-Ended User Query Dataset.
C.2 Dataset Statistics.
@@ Appendix C Open-Ended User Query Dataset
[…]
@@ Appendix E More Results
Figure 6: Performance Comparison across VBench Dimensions for Different Base Models. This visualization highlights the performance of all backbone models, including GPT-4o and Claude models, providing a comprehensive comparison in each dimension for different backbone models. Hatched portions indicate predictions within the exact range, and solid portions within an error margin of one range. Specific numerical results are provided in Table C.2 and Table 8
@@ E.1 Experiments on Different Base Models
Claude. We conducted validation experiments using the same setup as described in Section 4.1, but replaced the base model from GPT-4o to Claude-3.5-Sonnet, specifically using the claude-3-5-sonnet-20241022 version. Table 9 and Table 10 respectively present the time cost and the number of samples required for evaluations across VBench Huang et al. (2024a) and T2I-CompBench Huang et al. (2023) dimensions when the base model is replaced with Claude. Furthermore, Table  C.2 and Table 8 provide the evaluation accuracy results corresponding to these experiments. Notably, even with Claude as the base model, only a small number of samples and a few minutes are needed to achieve high-quality evaluation results. Figure 6 illustrates the evaluation outcomes using bar charts, comparing Claude and GPT as the base model.
@@
E.2 Evaluation Results of Open-Ended User Queries

### v3

@@ Abstract
Fan Zhang1∗, Shulin Tian2∗, Ziqi Huang2‡∗, Yu Qiao1🖂, Ziwei Liu2🖂 1Shanghai Artificial Intelligence Laboratory, 2S-Lab, Nanyang Technological University Equal Contributions. Project Lead. 🖂Corresponding Authors. zhangfan199903@163.com, {shulin002, ziqi002}@ntu.edu.sg https://vchitect.github.io/Evaluation-Agent-project/
footnotetext: Code is available
@@ 2.1 Visual Generation and Evaluation
Visual generative models have gained significant attention in recent years. However, unlike perception tasks, which have clear evaluation metrics such as accuracy, evaluating visual generative tasks is more challenging due to the absence of a definitive “ground truth” or single correct answer. Metrics such as FID Heusel et al. (2017) and FVD Unterthiner et al. (2018) are commonly used to measure the distance between generated samples and reference datasets. Recent benchmarks Huang et al. (2023, 2024a); Zheng et al. (2025); Huang et al. (2024b); Liu et al. (2023); Lee et al. (2023); Sun et al. (2024) provide multi-dimensional evaluations tailored to specific model capabilities. However, whether relying on metrics like FID Heusel et al. (2017) and FVD Unterthiner et al. (2018) or benchmarks like VBench Huang et al. (2024a), these evaluations are often time-intensive and limited in scope.
@@ 3.1 Preliminaries: Evaluation of Visual Generative Models
  Models
Consistency
Consistency
Smoothness
Degree
Quality
Quality
@@ 8 Ethical Considerations
9 Acknowledgment
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ A.2 Execution Stage
  Models
Consistency
Consistency
Smoothness
Degree
Quality
Quality
@@
Appendix B Experiment Implementation Details
For T2V tasks, we validate our evaluation approach by comparing the consistency of our Evaluation Agent with VBench’s original evaluation scheme across multiple dimensions, with respect to evaluation time, sample count, and final assessment results on four open-source T2V models. For VBench, we selected 15 evaluation dimensions, including Subject Consistency, Background Consistency, and Motion Smoothness, among others. Details could be referred to in Table 3.1. We used VBench’s original scheme to sample and evaluate four models on these dimensions, recording generation time, evaluation time, and results. To support a structured evaluation, we implemented a unified scoring pipeline aligned with VBench’s benchmark standards. Specifically, for each evaluation dimension, we first used the official metric tools to score a large number of video samples and partitioned the resulting sample-level scores into five predefined tiers: Very High, High, Moderate, Low, and Very Low, based on the score density distributions. For dimensions scored by proportions, such as Dynamic Degree, we constructed tiering thresholds based on results from $42$ models on VBench’s leaderboard. To ensure fairness, the Evaluation Agent was restricted to the tools and prompts provided by VBench. During evaluation, the agent answered targeted questions, invoked appropriate tools, and selected prompts accordingly. For each sampled video, it executed the relevant metric function to obtain a numerical score, which was then mapped into one of the five tiers using the predefined criteria. Over multiple rounds of evaluation, the agent aggregated these tiered results and determined the model’s overall performance level for each dimension. This process effectively translates intermediate quantitative observations into a final categorical assessment.
Appendix C Open-Ended User Query Dataset
C.1 Building an Open-Ended User Query Dataset.
C.2 Dataset Statistics.
@@ D.2 Agent in Action Modelling with Tool-Use
D.3 Recent Works on LLM-based Evaluation for Visual Generation
Recent works explore LLM-based agents for evaluating visual generative models. Qin et al. (2024) propose a question-answering agent that uses scene graphs to detect hallucinations in text-to-image diffusion models. Yang et al. (2025) introduce VideoGen-Eval, a benchmark for text-to-video generation that evaluates different aspects of video quality using LLM-guided reasoning. These studies reflect a growing interest in agent-based, interpretable evaluation frameworks for visual generation tasks.
@@ Appendix E More Results
Figure 6: Performance Comparison across VBench Dimensions for Different Base Models. This visualization highlights the performance of all backbone models, including GPT-4o and Claude models, providing a comprehensive comparison in each dimension for different backbone models. Hatched portions indicate predictions within the exact range, and solid portions within an error margin of one range. Specific numerical results are provided in Table A.3 and Table 8
@@ E.1 Experiments on Different Base Models
Claude. We conducted validation experiments using the same setup as described in Section 4.1, but replaced the base model from GPT-4o to Claude-3.5-Sonnet, specifically using the claude-3-5-sonnet-20241022 version. Table 9 and Table 10 respectively present the time cost and the number of samples required for evaluations across VBench Huang et al. (2024a) and T2I-CompBench Huang et al. (2023) dimensions when the base model is replaced with Claude. Furthermore, Table  A.3 and Table 8 provide the evaluation accuracy results corresponding to these experiments. Notably, even with Claude as the base model, only a small number of samples and a few minutes are needed to achieve high-quality evaluation results. Figure 6 illustrates the evaluation outcomes using bar charts, comparing Claude and GPT as the base model.
@@
E.2 Agent Reliability and Error Handling
While the Plan Agent is designed to reason in a structured, multi-step manner, agentic reasoning processes can still encounter failure modes, such as making unreasonable decisions, entering repetitive loops, or failing to converge on a clear evaluation. To mitigate these risks, our framework integrates several safeguards at both the agent and system levels. At the agent level, we adopt a ReAct-style prompting strategy, which encourages the Plan Agent to reason step-by-step and to explicitly articulate its thought process before each action. In particular, the agent is required to provide a justification (“thought”) before terminating the evaluation, helping ensure that such decisions are based on collected evidence rather than arbitrary heuristics. At the system level, we implement hard constraints such as a maximum number of evaluation rounds to prevent infinite loops and control evaluation cost. This has shown practical effectiveness, for example, as illustrated in Figure 8, where a repetitive loop triggered by Gemini was successfully halted by the framework protocol. Looking ahead, we believe it is valuable to further explore safeguard strategies for agent-based evaluation. While our current mechanisms offer basic robustness, improving fallback design and reliability remains an important direction for future work, especially as agent-based reasoning becomes more widely adopted.
E.3 Evaluation Results of Open-Ended User Queries